A stock just paid a dividend of $1.00. The dividend is expected to grow at 20.16% for three years and then grow at 3.34% thereafter. The required return on the stock is 13.82%. What is the value of the stock?

Answer format: Currency: Round to: 2 decimal places.

A stock just paid a dividend of $1.18. The dividend is expected to grow at 28.71% for five years and then grow at 3.19% thereafter. The required return on the stock is 13.03%. What is the value of the stock?

Answer format: Currency: Round to: 2 decimal places.

A stock just paid a dividend of $1.59. The dividend is expected to grow at 29.83% for two years and then grow at 4.24% thereafter. The required return on the stock is 10.17%. What is the value of the stock?

Answer format: Currency: Round to: 2 decimal places.

I would really appreciate the help!! :) Also it would be great if you could show the steps (not in excel) for better understanding~ Thank You
Answer #1

D0 = $1

The dividend paid in :

Year 1 : $1.2016 ($1 * 1.2016)

Year 2 : $1.4438

Year 3 : $1.7349

Year 4 : $1.7928 ( 1.7349 * 1.0334)

P3 = D4/ Re -g

= $1.7928  / 0.1382 - 0.0334

= $17.1069

The value of the stock is :

= $1.2016 / 1.1382 + $1.4438/1.1382^2 + $1.7349 /1.1382^3 + $17.1069 / 1.1382^3

= $14.9483

= $14.95 ( rounded off to two decimal places)

D0 = $1.18

D1 = $1.5188

D2 = $1.9548

D3= $2.516

D4= $3.2384

D5= $4.1681

D6 =$4.3011

P5 = D6/ Re - G

= $4.3011/ 0.1303 - 0.0319

= $43.7104

So, the value of the stock is : D1 + D2 + D3 + D4 + D5 + P5 ( DISCOUNTED AT 13.03%)

= $32.5523

= $32.55

The stock just paid a dividend of $1.59,

D1 = $1.59 *1.2983

= $2.0643

D2 = $2.6801

D3 = $2.6801 * 1.0424

= $2.7937

P3 = D3/ Re - g

= $2.7937/ 0.1017 - 0.0424

= $2.7937/ 0.0593

= $47.1113

The value of stock is : D1 + D2 + D3 +P3 (DISCOUNTED AT 10.17%)

= $41.5 ( rounded off to two decimal places)

= $23.35 ( rounded off to two decimal places)